OpenStudy (anonymous):

Which equation is the vertex form of y=-3x2+12x-7

OpenStudy (jdoe0001):

well, that's not in vertex form to begin with

OpenStudy (jdoe0001):

unless they're asking for the vertex form of that "standard form"

OpenStudy (jhannybean):

Vertex form is \[\large y=a(x-h)^2+k\]

OpenStudy (marissalovescats):

I think you will have to do a "complete the square" thing for this problem.

OpenStudy (marissalovescats):

She's supposed to put the formula into vertex form.

OpenStudy (marissalovescats):

I think you have to subtract 12 from the 7... I remember there being a rule about this because its like not on the side of an equal sign or something.

OpenStudy (marissalovescats):

I could be wrong but idk. -3*4 is -12 I'm unsure if I'm remembering correctly.

OpenStudy (jhannybean):

Nope. \[\large -3(x^2 +4x+4 -4) -7\]\[\large -3(x^2 +4x +4) -7+12\]

OpenStudy (jhannybean):

you multiply -4 by -3 and factor it out to complete the square.

OpenStudy (jhannybean):

You subtract -4 to balance the equation.

OpenStudy (jhannybean):

Small mistake.... \[\large y = -3x^2 +12x -7\]\[\large y=-3(x^2-4x)-7\]\[\large y= -3(x^2 -4x+4-4)-7\]\[\large y=-3(x-2)^2 -7+12\]\[\large y= -3(x-2)^2 +5\]